OpenStudy (elisaneedshelp):

For what acute angle θ is cos θ=sqrt3/2

OpenStudy (elisaneedshelp):

\[\cos \theta=\frac{ \sqrt{3} }{ 2 }\]

OpenStudy (anonymous):

Use cos inverse function in your calculator, or use the tables \[\cos ^{-1}\frac{ \sqrt{3} }{ 2 } = 30^{o}\]

OpenStudy (elisaneedshelp):

Thanks i didnt understand what it was asking so i thought it was something harder

OpenStudy (anonymous):

No problem :)